OpenStudy (anonymous):

find the antiderivative of (3x+10)(3x-10)

OpenStudy (anonymous):

we havent learned rules of multiplication when it comes to antiderivatives yet...so not sure how to handle it

OpenStudy (jusaquikie):

you have to expand it then you do it like the other ones. \[9x^2-100\]

OpenStudy (jusaquikie):

\[\frac{ 9x^3 }{ 3 }+100x +C\]

OpenStudy (anonymous):

oh ok so the antiderivative is 3x^3-100x

OpenStudy (anonymous):

ya plus the constant

OpenStudy (anonymous):

thx!

OpenStudy (jusaquikie):

-100 not +100 sorry

OpenStudy (anonymous):

i have another question about antiderivatives if you could help please :)

OpenStudy (jusaquikie):

sure

OpenStudy (anonymous):

the antiderivative of (x^.5+.5x^-.5)2 (the square root of x plus half times square root of x (in denominator))squared

OpenStudy (anonymous):

because if i do it the other way you did before (add the the expoenent) and divide by it.....it doesnt work because when you take the derivative you gotta take it of the inside function as well

OpenStudy (jusaquikie):

how bout if you distribute the ^2 ?

OpenStudy (jusaquikie):

x^2.5+.5x^1.5

OpenStudy (jusaquikie):

then you can do it like normal

OpenStudy (anonymous):

oooo ok thx

OpenStudy (anonymous):

wouldnt you multiply the 2 not add it .....so it would be x+.5x^-1

OpenStudy (jusaquikie):

nah think of 10^2 * 10^2 = 100*100=10,000 = 10^4

OpenStudy (jusaquikie):

or 10^4* 10^2 = 10,000 * 100 = 1,000,000 = 10^6

OpenStudy (anonymous):

but thats because they have the same base.....(a^x)^y = a^xy not a^(x+y)

OpenStudy (jusaquikie):

x^2 =(x)^2 = (x)x it's all the same

OpenStudy (jusaquikie):

pull out your calculator and try some examples just plug in any number for x

OpenStudy (anonymous):

3x^3-100x+constant